OpenStudy (anonymous):

Solve the equation. e^x-12e^-x-1=0

OpenStudy (anonymous):

e^x -12^(-x) -1 = 0

OpenStudy (anonymous):

\[e^x-12e ^{-x}-1=0 \] multiply by e^x \[e ^{2x}-12-e^x=0\] put \[e^x=t\] \[t^2-12-t=0\] find t and then x